OpenStudy (countrygirl1431):

If g (x)=2x + 1, find g (a + h)- g (a). 2h 2h + 2 h + 1

OpenStudy (countrygirl1431):

@needhelp247356

OpenStudy (anonymous):

k let me see

OpenStudy (anonymous):

only 3 options right?

OpenStudy (countrygirl1431):

yeah

OpenStudy (jhannybean):

Are you sure you're not looking for \(g(x+h)-g(x)\)?

OpenStudy (countrygirl1431):

all i know is the question

OpenStudy (jhannybean):

Unless \(x=a\) you cannot solve this function because the original function does not have an a.

OpenStudy (anonymous):

i got (d)

OpenStudy (jhannybean):

Let's see, now if we were solving for \(g(x+h)-g(x)\) it would go something like this: \[g(x+h) = 2(x+h)+1 = 2x+2h+1\]\[g(x) = 2x+1\]\[g(x+h)-g(x) = 2x+2h+1-(2x+1) = 2x+2h+1-2x-1 = \boxed{2h} \]

OpenStudy (countrygirl1431):

so 2h

OpenStudy (jhannybean):

Try revising your question, its not a, but should be x.

OpenStudy (countrygirl1431):

i cant revise it thats the question on my assignment

OpenStudy (jhannybean):

Ask your teacher if it's really g(a+h) - g(a) or if it's g(x+h)-g(x)

OpenStudy (countrygirl1431):

ok thanks

OpenStudy (jhannybean):

Np.